### v2

@@
Approaching Constrained Markov Decision Process
@@
Achieving $\tilde{O}(1/\varepsilon)$ Sample Complexity for Constrained Markov Decision Process
@@
We consider the reinforcement learning problem for the constrained Markov decision process (CMDP), which plays a central role in satisfying safety or resource constraints in sequential learning and decision-making. In this problem, we are given finite resources and a MDP with unknown transition probabilities. At each stage, we take an action, collecting a reward and consuming some resources, all assumed to be unknown and need to be learned over time. In this work, we take the first step towards deriving optimal problem-dependent guarantees for the CMDP problems. We derive a logarithmic regret bound, which translates into a $O(\frac{1}{\Delta\cdot\varepsilon}\cdot\log^{2}(1/\varepsilon))$ sample complexity bound, with $\Delta$ being a problem-dependent parameter, yet independent of $\varepsilon$. Our sample complexity bound improves upon the state-of-art $O(1/\varepsilon^{2})$ sample complexity for CMDP problems established in the previous literature, in terms of the dependency on $\varepsilon$. To achieve this advance, we develop a new framework for analyzing CMDP problems. To be specific, our algorithm operates in the primal space and we resolve the primal LP for the CMDP problem at each period in an online manner, with adaptive remaining resource capacities. The key elements of our algorithm are: i) a characterization of the instance hardness via LP basis, ii) an eliminating procedure that identifies one optimal basis of the primal LP, and; iii) a resolving procedure that is adaptive to the remaining resources and sticks to the characterized optimal basis.
@@ 1 Introduction
Reinforcement learning (RL) is pivotal in the realm of dynamic decision-making under uncertainty, where the objective is to maximize total reward through ongoing interaction with and learning from an enigmatic environment. Markov Decision Processes (MDPs) are a prevalent framework for encapsulating environmental dynamics. MDPs have been instrumental in various domains, such as video gaming (Mnih et al. 2013), robotics (Kober et al. 2013), recommender systems (Shani et al. 2005), and beyond. Yet, they fall short in accommodating additional constraints that may influence the formulation of the optimal policy and the decision-maker’s engagement with the uncertain environment. Often, in MDP applications, there are stringent constraints on utilities or costs, emanating from areas like safe autonomous driving (Fisac et al. 2018), robotics (Ono et al. 2015), and financial management (Tamar et al. 2012). These constraints might also symbolize limitations on resources within resource allocation contexts with applications to revenue management (Jia et al. 2023, Jiang 2023, Li et al. 2023). Constrained MDPs (CMDPs), as introduced in Altman (1999), enhance MDPs to factor in constraints affecting long-term policy results. In CMDPs, the decision-maker aims to optimize cumulative rewards while adhering to these constraints. Our paper focuses on CMDPs, and we aim to develop efficient algorithmic solutions.
@@ 1.1 Main Results and Contributions
We finally discuss the dependency of our sample complexity bound on problem parameters other than $\varepsilon$. We restrict to the MDP context without resource constraints. Denote by $\mathcal{S}$ the state set and $\mathcal{A}$ the action set. We show a sample complexity bound of $O\left(\frac{|\mathcal{S}|^{4}\cdot|\mathcal{A}|}{(1-\gamma)^{4}\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right)$, where $\Delta$ is the constant that represents the hardness of the underlying problem instance. Compared to the optimal worst-case sample complexity $O\left(\frac{|\mathcal{S}|\cdot|\mathcal{A}|}{(1-\gamma)^{3}\cdot\varepsilon^{2}}\right)$ that is achieved in a series of work (Sidford et al. 2018, Wainwright 2019, Wang 2020, Agarwal et al. 2020, He et al. 2021), our bound has a worse dependency over $|\mathcal{S}|$ and $1-\gamma$. This is due to our algorithm being LP-based and the dimension of the LP ($|\mathcal{S}|$ and $1-\gamma$) will influence our final bounds. However, our bound enjoys a better dependency in terms of $\varepsilon$. For the general CMDP problem, our bound will depend additionally on the conditional number of the constraint matrix in the LP formulation, which is a byproduct of the resolving LP heuristics (Vera and Banerjee 2021, Li et al. 2021c). However, our $\tilde{O}(1/\varepsilon)$ sample complexity bound depends polynomially on other parameters including $|\mathcal{S}|$, $|\mathcal{A}|$, $1-\gamma$, and the number of constraints.
@@ 1.2 Further Related Work
Deriving the (logarithmic) problem-dependent guarantees is a long-standing topic in online decision-making (Lai and Robbins 1985, Graves and Lai 1997). Notably, there has been a recent trend of deriving (logarithmic) problem-dependent guarantees for reinforcement learning problems, where a series of work (Zanette and Brunskill 2019, Simchowitz and Jamieson 2019, Mou et al. 2020, Foster et al. 2020, Dann et al. 2021, Wagenmaker et al. 2022, Wagenmaker and Foster 2023, Xu and Zeevi 2024, Duan and Wainwright 2024) has proposed various ways to characterize the hardness of the problem instance. A prevalent way to characterize the hardness is to define a constant that represents the policy sub-optimality gap between the optimal policy and a policy that takes a sub-optimal action given a state. Then, the logarithmic guarantees depend on this constant. Compared to the previous approaches and guarantees, our guarantees adopt this idea of policy sub-optimal gap. Moreover, our approach can handle resource or safety constraints that need to be satisfied by the policy. In this way, our approach generalizes to multi-objective (safe) RL problems. Safe reinforcement learning problems have been investigated extensively under various settings in a series of work (Liu et al. 2021a, Wei et al. 2021, Ding et al. 2021, Li et al. 2021b, Ghosh et al. 2022, Chen et al. 2022, Ding and Lavaei 2023, Germano et al. 2023, Gladin et al. 2023), where the $O(\sqrt{T})$ regret bounds or $O(1/\varepsilon^{2})$ sample complexity bounds are proved.
Our method is motivated by the online LP literature. The online LP problem (Molinaro and Ravi 2014, Agrawal et al. 2014, Gupta and Molinaro 2014) covers a wide range of applications through different ways of specifying the underlying LP, including secretary problem (Ferguson 1989), online knapsack problem (Arlotto and Xie 2020), quantity-based network revenue management problem (Jasin and Kumar 2012), network routing problem (Buchbinder and Naor 2009), matching problem (Mehta et al. 2007), etc. Notably, the problem has been studied under either (i) the stochastic input model where the coefficient in the objective function, together with the corresponding column in the constraint matrix is drawn from an unknown distribution $\mathcal{P}$, or (ii) the random permutation model where they arrive in a random order (Molinaro and Ravi 2014, Agrawal et al. 2014, Kesselheim et al. 2014, Gupta and Molinaro 2014). With an additional non-degeneracy assumption, the logarithmic regret has been derived for the quantity-based network revenue management problem (Jasin and Kumar 2012, Jasin 2014), the online LP problem (Li and Ye 2022), and the more general convex allocation problem (Ma et al. 2022). The non-degeneracy assumption has been relaxed in a series of work (Bumpensanti and Wang 2020, Vera and Banerjee 2021, Jiang et al. 2022, Ao et al. 2024) under various settings.
@@ 2 Preliminaries
There is a stochaastic reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}[0,1]$ and $K$ stochastic cost functions $c_{k}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}[0,1]$ for each $k\in[K]$. We also denote by $\hat{r}(s,a)=\mathbb{E}[r(s,a)]$ for each $(s,a)$ and $\hat{c}(s,a)=\mathbb{E}[c_{k}(s,a)]$ for each $(s,a)$. For any Markovian policy $\pi$, where the action of $\pi$ depends only on the current state and the action of $\pi$ is allowed to be randomized, we denote by $V_{r}(\pi,\mu_{1})$ the infinite horizon discounted reward of the policy $\pi$, with the formulation of $V_{r}(\pi,\mu_{1})$ given below:
@@ 7.2.2 Offline Learning
where $D$ is given in (49), $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal soluton to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 such that

### v3

@@
Sample Complexity for Constrained Markov Decision Process
@@
Achieving Instance-dependent Sample Complexity for Constrained Markov Decision Process
@@
We consider the reinforcement learning problem for the constrained Markov decision process (CMDP), which plays a central role in satisfying safety or resource constraints in sequential learning and decision-making. In this problem, we are given finite resources and a MDP with unknown transition probabilities. At each stage, we take an action, collecting a reward and consuming some resources, all assumed to be unknown and need to be learned over time. In this work, we take the first step towards deriving optimal problem-dependent guarantees for the CMDP problems. We derive a logarithmic regret bound, which translates into a $O(\frac{1}{\Delta\cdot\varepsilon}\cdot\log^{2}(1/\varepsilon))$ sample complexity bound, with $\Delta$ being a problem-dependent parameter, yet independent of $\varepsilon$. Our sample complexity bound is the first problem-dependent guarantee for CMDP problems. To achieve this advance, we develop a new framework for analyzing CMDP problems. To be specific, our algorithm operates in the primal space and we resolve the primal LP for the CMDP problem at each period in an online manner, with adaptive remaining resource capacities. The key elements of our algorithm are: i) a characterization of the instance hardness via LP basis, ii) an eliminating procedure that identifies one optimal basis of the primal LP, and; iii) a resolving procedure that is adaptive to the remaining resources and sticks to the characterized optimal basis.
@@ 1 Introduction
Reinforcement learning (RL) is pivotal in the realm of dynamic decision-making under uncertainty, where the objective is to maximize total reward through ongoing interaction with and learning from an enigmatic environment. Markov Decision Processes (MDPs) are a prevalent framework for encapsulating environmental dynamics. MDPs have been instrumental in various domains, such as inventory management (Alvo et al. 2023), video gaming (Mnih et al. 2013), robotics (Kober et al. 2013), recommender systems (Shani et al. 2005), and beyond. Yet, they fall short in accommodating additional constraints that may influence the formulation of the optimal policy and the decision-maker’s engagement with the uncertain environment. Often, in MDP applications, there are stringent constraints on utilities or costs, emanating from areas like safe autonomous driving (Fisac et al. 2018), robotics (Ono et al. 2015), and financial management (Tamar et al. 2012). These constraints might also symbolize limitations on resources within resource allocation contexts with applications to revenue management (Jia et al. 2023, Jiang 2023, Li et al. 2023), inventory management and queueing systems (Chen et al. 2021). Constrained MDPs (CMDPs), as introduced in Altman (1999), enhance MDPs to factor in constraints affecting long-term policy results. In CMDPs, the decision-maker aims to optimize cumulative rewards while adhering to these constraints. Our paper focuses on CMDPs, and we aim to develop efficient algorithmic solutions.
@@ 1.1 Main Results and Contributions
We finally discuss the dependency of our sample complexity bound on problem parameters other than $\varepsilon$. We restrict to the MDP context without resource constraints. Denote by $\mathcal{S}$ the state set and $\mathcal{A}$ the action set. We show a sample complexity bound of $O\left(\frac{|\mathcal{S}|^{4}\cdot|\mathcal{A}|}{(1-\gamma)^{4}\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right)$, where $\Delta$ is the constant that represents the hardness of the underlying problem instance. Compared to the optimal worst-case sample complexity $O\left(\frac{|\mathcal{S}|\cdot|\mathcal{A}|}{(1-\gamma)^{3}\cdot\varepsilon^{2}}\right)$ that is achieved in a series of work (Sidford et al. 2018, Wainwright 2019, Wang 2020, Agarwal et al. 2020, He et al. 2021), our bound has a worse dependency over $|\mathcal{S}|$ and $1-\gamma$. This is due to our algorithm being LP-based and the dimension of the LP ($|\mathcal{S}|$ and $1-\gamma$) will influence our final bounds. However, our bound enjoys a better dependency in terms of $\varepsilon$. For the general CMDP problem, our bound will depend additionally on the conditional number of the constraint matrix in the LP formulation, which is a byproduct of the resolving LP heuristics (Vera and Banerjee 2021, Li et al. 2021c). However, our $\tilde{O}(1/\varepsilon)$ sample complexity bound depends polynomially on other parameters including $|\mathcal{S}|$, $|\mathcal{A}|$, $1-\gamma$, and the number of constraints. Partial results of this paper have been accepted at NeurIPS 2024 conference. Compared with the conference version, the current paper not only enriches the interpretations significantly and includes complete detailed proofs, but also includes the following key extensions.
@@ 1.2 Further Related Work
Deriving the (logarithmic) problem-dependent guarantees is a long-standing topic in online decision-making (Lai and Robbins 1985, Graves and Lai 1997). Notably, there has been a recent trend of deriving (logarithmic) problem-dependent guarantees for reinforcement learning problems, where a series of work (Zanette and Brunskill 2019, Simchowitz and Jamieson 2019, Mou et al. 2020, Foster et al. 2020, Dann et al. 2021, Bhandari and Russo 2021, Wagenmaker et al. 2022, Wagenmaker and Foster 2023, Xu and Zeevi 2024, Duan and Wainwright 2024, Chen et al. 2024) has proposed various ways to characterize the hardness of the problem instance. A prevalent way to characterize the hardness is to define a constant that represents the policy sub-optimality gap between the optimal policy and a policy that takes a sub-optimal action given a state. Then, the logarithmic guarantees depend on this constant. Compared to the previous approaches and guarantees, our guarantees adopt this idea of policy sub-optimal gap. Moreover, our approach can handle resource or safety constraints that need to be satisfied by the policy. In this way, our approach generalizes to multi-objective (safe) RL problems. Safe reinforcement learning problems have been investigated extensively under various settings in a series of work (Liu et al. 2021a, Wei et al. 2021, Ding et al. 2021, Li et al. 2021b, Ghosh et al. 2022, Chen et al. 2022, Ding and Lavaei 2023, Germano et al. 2023, Gladin et al. 2023), where the $O(\sqrt{T})$ regret bounds or $O(1/\varepsilon^{2})$ sample complexity bounds are proved.
Our method is motivated by the online LP literature. The online LP problem (Molinaro and Ravi 2014, Agrawal et al. 2014, Gupta and Molinaro 2014) covers a wide range of applications through different ways of specifying the underlying LP, including secretary problem (Ferguson 1989), online knapsack problem (Arlotto and Xie 2020), quantity-based network revenue management problem (Jasin and Kumar 2012), network routing problem (Buchbinder and Naor 2009), matching problem (Mehta et al. 2007), etc. Notably, the problem has been studied under either (i) the stochastic input model where the coefficient in the objective function, together with the corresponding column in the constraint matrix is drawn from an unknown distribution $\mathcal{P}$, or (ii) the random permutation model where they arrive in a random order (Molinaro and Ravi 2014, Agrawal et al. 2014, Kesselheim et al. 2014, Gupta and Molinaro 2014). With an additional non-degeneracy assumption, the logarithmic regret has been derived for the quantity-based network revenue management problem (Jasin and Kumar 2012, Jasin 2014), the online LP problem (Li and Ye 2022), and the more general convex allocation problem (Ma et al. 2022). The non-degeneracy assumption has been relaxed in a series of work (Bumpensanti and Wang 2020, Vera and Banerjee 2021, Jiang et al. 2022, Wei et al. 2023, Ao et al. 2024, Banerjee and Freund 2024) under various settings.
@@ 2 Preliminaries
There is a stochaastic reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}[0,1]$ and $K$ stochastic cost functions $c_{k}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}[0,1]$ for each $k\in[K]$. We also denote by $\hat{r}(s,a)=\mathbb{E}[r(s,a)]$ for each $(s,a)$ and $\hat{c}_{k}(s,a)=\mathbb{E}[c_{k}(s,a)]$ for each $(s,a)$, for each $k$. For any Markovian policy $\pi$, where the action of $\pi$ depends only on the current state and the action of $\pi$ is allowed to be randomized, we denote by $V_{r}(\pi,\mu_{1})$ the infinite horizon discounted reward of the policy $\pi$, with the formulation of $V_{r}(\pi,\mu_{1})$ given below:
@@ 7.2.2 Offline Learning
where $D$ is given in (49), $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal solution to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 such that
@@
7.3 Numerical Studies
We implement our Algorithm 2 to study the numerical performance. We consider a CMDP problem with the state space $|\mathcal{S}|=10$ and the action space $|\mathcal{A}|=10$. We set the discount factor $\gamma=0.7$. We then randomly generate the probability transition kernel $P$. To be specific, for each state $s\in\mathcal{S}$, action $a\in\mathcal{A}$, and the future state $s^{\prime}\in\mathcal{S}$, we uniformly generate a randomly variable $p_{s,a,s^{\prime}}$. Then, the transition probability is defined as $P(s^{\prime}|s,a)=\frac{p_{s,a,s^{\prime}}}{\sum_{s^{\prime\prime}\in\mathcal{S}}p_{s,a,s^{\prime\prime}}}$. For each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected reward $\hat{r}(s,a)$ is uniformly generated from the interval $[1,2]$ (with the reward for the first action set to be 00). The actual reward $r(s,a)=\hat{r}(s,a)+\eta$, where $\eta$ is uniformly distributed among $[-0.5,0.5]$. There are $K=5$ constraints and for each constraint $k\in[K]$ and each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the expected cost $\hat{c}_{k}(s,a)$ to be uniformly generated from $[1,2]$. The actual cost $c_{k}(s,a)=\hat{c}_{k}(s,a)+\eta^{\prime}$, where $\eta^{\prime}$ is uniformly distributed among $[-0.5,0.5]$.
For each total iterations $N$, We apply Algorithm 2 and obtain the output $\bm{q}^{1},\dots,\bm{q}^{N}$. We compare $\bar{\bm{q}}^{N}$ with the optimal occupancy measure. Since our algorithm is a randomized algorithm, we study the performance of our algorithm in expectation. To be specific, given the problem instance and a fixed $N$, we implement our algorithm repeatedly for $M=500$ rounds. Denote by $\bar{\bm{q}}^{N}_{m}$ the output of our Algorithm 2 at round $m$, for $m\in[M]$. We define the error term as $\text{Err}(N)=\frac{1}{M}\cdot\sum_{m=1}^{M}\|\bar{\bm{q}}_{m}^{N}-\bm{q}^{*}\|_{1}/\|\bm{q}^{*}\|_{1}$. We study how the error term $\text{Err}(N)$ scales with $N$. The results are displayed in Figure 2. As we can see, the error term $\text{Err}(N)$ decreases quickly with respect to $N$. Moreover, since our Algorithm 2 requires only solving a set of linear equations in each iteration, the computation cost of our Algorithm 2 is also moderate.
Figure 2: The computational performance of Algorithm 2. The x-label denotes the size of $N$, while the y-label denotes the error term $\text{Err}(N)$.